You are here: Foswiki>Commissioning Web>Com20140124 (25 Jan 2014, TarasGolota)Edit Attach

Technical Observing - 20140124 UT

Observers: DThompson (LBTO), JHill (Tucson - remote room) DMiller (home under the weather)
Telescope Operator: GBechetti (LBTO)
AO Support:: DMiller (Tucson)
Telescope Support:: JUrban (LBTO)
Instrument Support:: ESolheid (LBTO)
Engineering Support: DAshby (Tucson)

Summary

This technical night is a mix of E-time (TCS tests and IQ observations and instrument configuration tests) and D-time (staff science and AZ science).

Rotational Offsets for Issue 5018

Preference Offloading for Issue 5030

Validation of Patches to PCS, GCS and

Details

Open

We are using PCS patch PCSV6.27c. Preference offloading is now enabled by default in binocular mode.

The Z6 LUT on the primary is turned off for this initial observation.

Also using GCS v2.14.14 for this initial observation.

monocular MODS1 Science Program with DThompson and FBian

01:05 ACQUIRE preset as MODS1 on Left Only - found M5 star at IE=-35 CA=-55 (off at the edge of the MODS1 acq field), but it was a binary w/o good coordinates. Open-loop collimation was good, just a bit of defocus - image was ~1.2 arcsec.

01:14 ACTIVE preset, but waiting 4 minutes for wfs to get dark enough. Sky has a fair bit of scattered cirrus. Seeing about 1 arcsec on guider. Temperature is near freezing.

01:30 Guiding residuals are 30% larger in Y (0.09 rms) than X (0.07 rms). Is this the same effect Christian et al saw last week? Vignetting effects on wfs give the guide star a slight tail to the right (X).

01:38 starting the science exposure

01:48 at EL=67.8 SX M1 X=0.046 Y=0.030 RX=-1.7 RY=20.7
SX M2 X=-3.116 Y=+1.425 Z=1.978 RX=242.4 Y=-153.5

02:28 Guiding residuals are now equal in both directions (0.08 rms).

02:35 Science finished in thickening clouds

TCS Testing - Offsets in rotation

Still using MODS1 on LDG ....

02:38 offset 0 0 ROT=-30 DETXY REL (probe was at -37 -133 SFP and moved to 47 -131) The guide star was found within a fraction of an arcsec of the desired position. More tests below......

02:42 Reconfigure to LFBG / RFBG and to deal with LUCI1 temperature issues. Instrument electronics cooling problems?

02:48 Stop and restart both GCSs to implement patch v2.14.16.

02:54 Reconfigure complete, but still at zenith for LUCI1 issues.

03:08 LUCI1 issues controlled.

Authorize as LUCI on both sides.

03:11 one stowpin didn't retract.......

03:12 ACQUIRE SYNC presets on both sides .... Star on left has significant coma (fan is ~4 arcsec across), and we didn't find star on the right. Right collimation was OK after spiraling.

03:27 moving to a brighter star

SX IE=-50 CA=-85 DX IE=+8 CA=-201

03:42 Home SX M3, but that did nothing to change the position of the star.

PCS demands are SX Tip -11.5 DX Tip +14.5 SX Tilt +72.3 DX Tilt -55.33

03:53 Authorize as LUCI@LFBG only

03:56 ACQUIRE preset on left - pointing and bad collimation are the same.

03:57 Init SX M2 hexapod - pointing and bad collimation are the same.

04:00 Reset and reset SX Shell - pointing and bad collimation are the same.

04:02 Authorize as LUCI on both sides

04:06 ACTIVE SYNC preset to BS9107

04:10 "Initialize" EL LUT on SX M2. This changes X position by -1.5 mm, which then makes a big change in preference. ???????

04:17 ACQUIRE SYNC presets back to bright star (star not found on left, star well centered at previous IE/CA on right) After some spiraling, we find the SX star which is no longer comatic (only a bit of focus)!

Field Aberration Data at RFBG

04:22 Active preset on BS9107

Hmmm, maybe Geno/John had some bad global offsets in his PSF M2 gui. That could explain our collimation issues. But did he send any focus to SX M2 prior to 4 UT?

Now we don't have any spots on the SX wfs ......

04:28 Active preset on SX async

04:30 Stop and restart GCSL. Preset again.

Now we have the proper "Image didn't arrive within timeout.

04:35 Stop/Start AGw1 (and GCS again)

Data lives in /home/lbto/data/20140124_rfbg

04:39 field_collect, start=1, limit=300 on right with BS9107_new.list

Aha! on agw1-cam we see this message: Warning Dec 11, 2013 11:21:22 PM Controller 0 Unclean shutdown detected: unit=0
But we were getting guider images? And Doug/Taras were using it this afternoon.

NOTE by T.G (20140124): These 2 lines above are irrelevant (old and they do not affect camera operation according to M.L.) these errors indicate possible HW issue with HDD.

04:46 warm boot agw1-cam

2014-01-24T04:50:21.190493+00:00 tcs4 LBT_left_GCS: AzCam: bad reply from server while trying to set ROI
2014-01-24T04:50:21.190500+00:00 tcs4 LBT_left_GCS: AzCam cmd was: SetROI 16 536 40 508 2 2
2014-01-24T04:50:21.190506+00:00 tcs4 LBT_left_GCS: AzCam reply was: 
2014-01-24T04:50:21.190511+00:00 tcs4 LBT_left_GCS: AzCam: trying to resend command (3/3)
2014-01-24T04:50:23.190602+00:00 tcs4 LBT_left_GCS: AzCam: send('SetROI 16 536 40 508 2 2'): mutex for sending command to AzCam timed out, possible deadlock, giving up!
2014-01-24T04:50:23.190624+00:00 tcs4 LBT_left_GCS: AzCam: bad reply from server while trying to set ROI
2014-01-24T04:50:23.190631+00:00 tcs4 LBT_left_GCS: AzCam cmd was: SetROI 16 536 40 508 2 2
2014-01-24T04:50:23.190636+00:00 tcs4 LBT_left_GCS: AzCam reply was: 
2014-01-24T04:50:23.190641+00:00 tcs4 LBT_left_GCS: AzCam: trying to resend command (4/3)
2014-01-24T04:50:25.190743+00:00 tcs4 LBT_left_GCS: AzCam: send('SetROI 16 536 40 508 2 2'): mutex for sending command to AzCam timed out, possible deadlock, giving up!
2014-01-24T04:50:25.190765+00:00 tcs4 LBT_left_GCS: AzCam: bad reply from server while trying to set ROI
2014-01-24T04:50:25.190772+00:00 tcs4 LBT_left_GCS: AzCam cmd was: SetROI 16 536 40 508 2 2
2014-01-24T04:50:25.190778+00:00 tcs4 LBT_left_GCS: AzCam reply was: 
2014-01-24T04:50:25.190783+00:00 tcs4 LBT_left_GCS: AzCam: trying to resend command (5/3)
2014-01-24T04:50:25.195854+00:00 tcs4 LBT_left_GCS: PROFILING: getAcquisitionImage: requesting full CCD readout ... <===
2014-01-24T04:50:25.195869+00:00 tcs4 LBT_left_GCS: size: full CCD, exptime: 2000 ms

NOTE by T.G (20140124): According to log command SetROI 16 536 40 508 2 2 was not executed by AzCam. IT# 5033.

05:34 Stop field_collect

05:35 reboot azcamserver from the button, restart GCSs, but guidecam init failed

05:40 Stop GCSs, Stop AGW1, reboot azcamserver, reboot agw1-cam ,,,, and that finally worked.

Results from Doug: DX TIP +35 arcsec dIE -35 arcsec

DX M2 Manual X=0.000 Y=-1.787 Z=0 RX=343.7 RY=0

05:50 RIGHT field_collect, start=1, limit=300

05:55 LEFT field_collect, start=1, limit=300

===> Error in wfsc_read_info: file not found
       /Repository/GCSfiles/dailyimages/left_wfscimage000131.info
         => Could not process /Repository/GCSfiles/dailyimages//left_wfscimage000131.fits
              Probably could not find info file
              due to bad WFSC image

Because GCS was using different numbers on the subimage -- See Issue 5031.

-rw-rw-r--. 1 telescope domain users   1772 Jan 24 06:19 /Repository/GCSfiles/dailyimages/left_wfscimage000203.info
-rw-rw-r--. 1 telescope domain users 558720 Jan 24 06:19 /Repository/GCSfiles/dailyimages/left_wfscimage000203.fits
-rw-rw-r--. 1 telescope domain users  77760 Jan 24 06:19 /Repository/GCSfiles/dailyimages/left_wfscsubimage000158.fits
-rw-rw-r--. 1 telescope domain users   1766 Jan 24 06:19 /Repository/GCSfiles/dailyimages/left_wfscimage000204.info
-rw-rw-r--. 1 telescope domain users 558720 Jan 24 06:19 /Repository/GCSfiles/dailyimages/left_wfscimage000204.fits
-rw-rw-r--. 1 telescope domain users  77760 Jan 24 06:20 /Repository/GCSfiles/dailyimages/left_wfscsubimage000159.fits
-rw-rw-r--. 1 telescope domain users   1760 Jan 24 06:20 /Repository/GCSfiles/dailyimages/left_wfscimage000205.info
-rw-rw-r--. 1 telescope domain users 558720 Jan 24 06:20 /Repository/GCSfiles/dailyimages/left_wfscimage000205.fits
-rw-rw-r--. 1 telescope domain users  77760 Jan 24 06:21 /Repository/GCSfiles/dailyimages/left_wfscsubimage000160.fits
-rw-rw-r--. 1 telescope domain users   1777 Jan 24 06:21 /Repository/GCSfiles/dailyimages/left_wfscimage000206.info
-rw-rw-r--. 1 telescope domain users 558720 Jan 24 06:21 /Repository/GCSfiles/dailyimages/left_wfscimage000206.fits

06:20 Stop and restart GCSL to fix above - didn't help

Go to tcs4 and make a fake subimage file in GCS directory, but GCS is smarter than I am - that didn't work either. Doug created Issue 5031.

06:30 LEFT field_collect ........ failed. We used up too much time with Issue 5031, so we did not get any confirming field aberration data on LFBG tonight.

06:58 RIGHT field_collect is finished. This data set says that our first adjustment was slightly too large. Perhaps it should have been 25 arcsec rather than 35. There are issue of extra noise in the Z6 plot. Doug will study the data more carefully in the daytime. The TIP-TILT CORRECTIONS TO DX M2 tonight were NOT permanently applied to the collimation models.

Fast (but not too fast) Guide and WFS data

07:02 SYNC ACTIVE presets to AO264 at EL=75

IS The DETSEC MATH WRONG in the wfsc headers? CHECK THIS IN DAYTIME ---> DETSEC versus CCDSEC

07:21 selectAGW right LUCI_R.cfg

07:26 200 fast wfs and guide images on right side. It is taking 3.75 sec for 1.0 sec exposure images. Perhaps slower because guide and wfs are running together.

07:34 selectAGW left LUCI_L.cfg

07:40 200 fast guider only images on right side 500 ms images are taking 1.2 sec.

Doug took left side data in /tmp on tcs4, but that wasn't much faster. He copied them to: XXXXXXXX

WFS Camera problems

07:47 Restore the GCS config files to the normal ones, and selectAGW on both sides (i.e. selectAGW right LUCI_R.cfg).

Guide images are OK, WFS images are goofy -- did selectAGW not intialize the controllers properly? Same problem we had earlier.

TCS Testing - Offsets in rotation

07:55 Off-axis preset to gs=6 at PA~180 on both sides

07:56 RIGHT offset 0 0 ROT=-30 DETXY REL --- missed by about 4 arcsec on the first guide image. Pointing origin base was 0, 0 for this test (not the earlier one).

08:00 LEFT offset 0 0 ROT=-30 DETXY REL (finished at 08:00:27 I didn't see the guide image go by) Looking at the guide plot shows there was no problem.

This rotation offset error seems to only happen at RFBG. Issue 5018 has been updated.

QSO Spectroscopy with MODS1

08:05 Reconfigure for MODS1

Turn the rest of the night over to Olga for QSO spectroscopy

-- JohnHill - 24 Jan 2014
Topic revision: r4 - 25 Jan 2014, TarasGolota
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback